OpenStudy (toxicsugar22):

9y^2-8x^2+72y+16x+64=0 give the exact coordinates of any four points on the hyperbola

myininaya (myininaya):

let x=0 solve for y let y=0 solve for x you should get two solutions for each time so that is 4 solutions

OpenStudy (toxicsugar22):

can u help me put x for 0

myininaya (myininaya):

x=0 so we have 9y^2+72y+64=0 do you know how to solve this?

OpenStudy (toxicsugar22):

YES

OpenStudy (toxicsugar22):

ITS A quadractic equation

myininaya (myininaya):

thats right you sound happy lol i think

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

cant we foil

OpenStudy (toxicsugar22):

to make it smaller numbers

myininaya (myininaya):

i dont think there are any common factors but you see if the polynomial is factorable if you like

OpenStudy (anonymous):

hey myininaya can you help me out after with my trapezoidal rule question:)

OpenStudy (toxicsugar22):

so how do i do this

OpenStudy (toxicsugar22):

u do it and we will compare our answer

OpenStudy (toxicsugar22):

ok i got mine

OpenStudy (toxicsugar22):

answer

OpenStudy (toxicsugar22):

did u get an anser

OpenStudy (anonymous):

this is weird i never recall ever learning hyperbola's in high school or college.

OpenStudy (toxicsugar22):

myininama are you still there

myininaya (myininaya):

sorry i was away

myininaya (myininaya):

\[y=\frac{-72\pm \sqrt{72^2-4*9*64}}{2*9}\]

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

i got that

OpenStudy (toxicsugar22):

and then help me furthur

myininaya (myininaya):

with simplifying?

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

please

myininaya (myininaya):

the bottom is 18 because 2*9=18 i would be 72^2-4*9*64 in my calculator if i were you

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

exact answer i get is \[y=(-72\pm \sqrt{4968})\div18\]

myininaya (myininaya):

y=(-72+ sqrt{329472})/18 y=(-72- sqrt{329472})/18

OpenStudy (toxicsugar22):

that is te answer

myininaya (myininaya):

lol hes right i accidently put + between the b^2 and the 4ac

myininaya (myininaya):

hey beehive if you want your fractions to look cooler you can use frac{5}{7} and this will give you 5/7 in the equation editor if you want

OpenStudy (anonymous):

ok thanks.

OpenStudy (toxicsugar22):

so myininaya how do u go from there

myininaya (myininaya):

what do you mean?

myininaya (myininaya):

you can find the other two points now if you want

OpenStudy (anonymous):

\[y={-4} \pm \frac{\sqrt{4968}}{18}\]

myininaya (myininaya):

we already let x be 0 now lets let y be 0 and we will stil get a quadratic like -8x^2+16x+64=0

myininaya (myininaya):

divide both sides by -8 so we have x^2-2x-8=0 (x-4)(x+2)=0 x=4 and x=-2 so our four coordinates are (4,0), (-2,0), (0,-4+sqrt(4968)/18),(0,-4-sqrt(4968)/18)

OpenStudy (toxicsugar22):

i dont think the last two points are correct

OpenStudy (toxicsugar22):

can we simply it

myininaya (myininaya):

what did you get when u did 72^2-4(9)(36) i dont have my calculator if it is 4968 then those are corret above

myininaya (myininaya):

i only have this calcutor on computer and i dont like to use it

OpenStudy (toxicsugar22):

i got 3888

OpenStudy (toxicsugar22):

inm i got 2880

myininaya (myininaya):

72^2=5184 4*9*64=2304 i mean 64 not 36

OpenStudy (anonymous):

myininaya where did you get 36. that's what i mean.

myininaya (myininaya):

2880

OpenStudy (anonymous):

2880 here also

myininaya (myininaya):

lol that should be under the square root sign not 4 thousand something number

OpenStudy (toxicsugar22):

ok thanks i got it

OpenStudy (anonymous):

squareroot of 4968 i think

myininaya (myininaya):

but 72^2-4(9)(36)=2880

myininaya (myininaya):

i mean 64 not 36